Grade 11Algebra

bag contain 5 apples and 7 oranges and another bag contain 4 apples and 8 oranges. A fruit picked out from each bag. Find the probability that the fruits are both apples or both oranges.

Sakkshi
8 Years agoGrade 11

1 Answer

Arun
8 Years ago
Dear Sakkshi
 
P(both oranges) = (7C1/12C1) * (8C1/12C1) = (7/12) * (8/12)
 = 56/144
 
P(both apples) = (5C1/12C1) * (4C1/12C1)
 = (5/12) * (4/12)
 = 20/144
 
Hence
 
P(both apples or both oranges) = (56 + 20)/144 
 = 76/144
 
 
 
Regards
Arun (askIITians forum expert)